OpenStudy (anonymous):

what is 6/i in standard form?

OpenStudy (anonymous):

\[\frac{6}{i}*\frac{i}{i}=\frac{6i}{i^2}=\frac{6i}{-1}=-6i\]

OpenStudy (anonymous):

wait why multiply by i/i?

OpenStudy (anonymous):

easier is to memorize this relationship\[-i*i=-1\]

OpenStudy (anonymous):

is that the conjugate?

OpenStudy (anonymous):

you can't have i in the denominator because it is a radical by definition\[i=\sqrt{-1}\]so you are in effect rationalizing the denominator

OpenStudy (anonymous):

k i think i get it

jimthompson5910 (jim_thompson5910):

you want to make the denominator a real number, so you multiply i by itself to make it i^2 or -1 (you must do this to both numerator and denominator)

OpenStudy (anonymous):

a-bi is the conjugate of a+bi (the complex conjugate is the imaginary part of the complex number negated)

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

welcome